### v1

# Advancing Multi-Party Dialogue Systems with Speaker-ware Contrastive Learning

Abstract
Dialogue response generation has made significant progress, but most research has focused on dyadic dialogue. In contrast, multi-party dialogues involve more participants, each potentially discussing different topics, making the task more complex. Current methods often rely on graph neural networks to model dialogue context, which helps capture the structural dynamics of multi-party conversations. However, these methods are heavily dependent on intricate graph structures and dataset annotations, and they often overlook the distinct speaking styles of participants. To address these challenges, we propose CMR, a Contrastive learning-based Multi-party dialogue Response generation model. CMR uses self-supervised contrastive learning to better distinguish "who says what." Additionally, by comparing speakers within the same conversation, the model captures differences in speaking styles and thematic transitions. To the best of our knowledge, this is the first approach to apply contrastive learning in multi-party dialogue generation. Experimental results show that CMR significantly outperforms state-of-the-art models in multi-party dialogue response tasks.
Advancing Multi-Party Dialogue Systems with Speaker-ware Contrastive Learning
Zhongtian Hu1, Yiwen Cui1, Ronghan Li2, Meng Zhao3, Lifang Wang1, 1School of Computer Science and Engineering, Northwestern Polytechnical University, 2School of Computer Science and Technology, Xidian University, 3School of Artificial Intelligence and Big Data, Henan University of Technology ahxchzt@mail.nwpu.edu.cn, wanglf@nwpu.edu.cn
1 Introduction
In recent years, dialogue response generation has made notable progress Ide and Kawahara (2021); Liu et al. (2021c), largely due to advancements in language models. These improvements have enhanced the generation of coherent and contextually relevant responses in applications such as customer service, virtual assistants, and conversational agents Hu et al. (2024); Wu et al. (2022); Komeili et al. (2022). However, most research has focused on dyadic dialogues Cai and Cai (2022); Sun et al. (2021); Chen et al. (2022), involving only two participants. In contrast, multi-party dialogues, common in real-world scenarios like group discussions, meetings, and social media, present unique structural challenges. As shown in Figure 1, while dyadic dialogues follow natural speaker rotations, multi-party dialogues require careful management of information flow between multiple participants Gu et al. (2022b); Ganesh et al. (2023), posing significant challenges for response generation models.
Figure 1: This figure illustrates a multi-party dialogue involving four participants: John, Mike, Sarah, and Emma. Each participant engages in discussions on different topics, demonstrating the dynamics and complexity of multi-party conversations.
To address these complexities, many approaches utilize graph neural networks (GNNs) to model dialogue context Hu et al. (2019); Wang et al. (2020); Shi and Huang (2019). For example, Gu et al. (2022a) proposed the HeterMPC model, which uses utterance and interlocutor nodes with six meta relations to construct a heterogeneous graph that captures dialogue interactions. Similarly, Gu et al. (2023) introduced a fully connected graph to model utterance-interlocutor relationships, incorporating latent addressee inference. While GNNs effectively model the dependencies in multi-party dialogues, they have drawbacks. The complexity of graph structures and the need for extensive annotations, such as addressee labels Li and Zhao (2023), make these methods computationally expensive and challenging to scale in real-world applications.
Given these challenges, we propose CMR (Contrastive Learning-based Multi-Party Response Generation). Our approach uses contrastive learning to understand the structure and dynamics of multi-party dialogues without relying on complex graph structures or extensive annotations. The model follows a two-stage training process. In the first stage, only the encoder is trained. Through contrastive learning, the model contrasts utterances from different speakers within the same conversation and those from random speakers across the dataset. This helps the model distinguish between different facts and speaking styles, enhancing speaker awareness and leading to more relevant responses. In the second stage, the model is jointly trained on a response generation task with contrastive learning. Here, the model contrasts the dialogue context with generated responses (via beam search), the gold response, the target speaker’s previous utterance, and other speakers’ responses within the same conversation. This stage improves response quality by enhancing the model’s understanding of conversation themes.
We validated our approach using the Friends dataset. Both automatic evaluation results and human evaluation results demonstrate that our model significantly outperforms the baseline models.
Our primary contributions are as follows:
  • We propose CMR, a novel multi-party dialogue response generation model that leverages contrastive learning to better understand the structure and dynamics of multi-party dialogues. To the best of our knowledge, this is the first work on multi-party dialogue response generation based on contrastive learning.
  • We introduce a two-stage training process: the first stage focuses on distinguishing the speaking styles and facts among different participants through contrastive learning, and the second stage jointly trains the model with the response generation task and contrastive learning to improve the quality of responses and the model’s understanding of dialogue themes.
  • Both automatic evaluation results and human evaluation results on the Friends dataset demonstrate that our model significantly outperforms baseline models and achieves SOTA.
2 Related Work
Research on multi-party dialogue response generation is relatively sparse, but there are some pioneering works in this field. Additionally, we review some methods that combine contrastive learning with generation tasks.
2.1 Graph Neural Networks for Multi-Party Dialogue
Several studies have explored the use of GNNs to capture the intricate relationships in multi-party dialogues. For instance, GSN (Graph-Structured Network) introduced by Hu et al. (2019). models multi-party dialogues as graphs rather than sequences. GSN employs an utterance-level graph-structured encoder to model information flow along the graph and incorporates speaker information to improve performance Similarly, Gu et al. (2022a) proposed the HeterMPC model, which employs heterogeneous graph networks to model interactions between interlocutors and utterances using multiple types of nodes and edges. Furthermore, another approach by Gu et al. (2023) constructed fully connected graphs to model relationships between all utterances and interlocutors in a conversation. While addressee labels are not entirely unnecessary, their model is designed to infer latent addressees through a graph prediction model integrated into a Variational Inference framework, making it robust even when some addressee information is missing.
2.2 Enhanced Speaker-Aware Approaches
Ma et al. (2023) proposed an Enhanced Speaker-Aware Multi-Party Multi-Turn Dialogue Comprehension model that integrates speaker-related features at both the token and discourse levels. By applying a speaker mask for attention and using heterogeneous graph networks for comprehensive speaker-aware discourse clues, this model achieves state-of-the-art performance on datasets such as MolweniLi et al. (2020) and FriendsQAYang and Choi (2019). Liu et al. (2021b) proposed an Utterance-Aware and Speaker-Aware Multi-turn Dialogue Representation model (MDFN) that integrates utterance-level and speaker-level features using a mask attention mechanism. The MDFN achieve amazing performance on several Datasets for response selection task. These methods highlights the importance of incorporating speaker-aware signals to improve dialogue comprehension and response generation.
2.3 Contrastive Learning for Text Generation
Contrastive learning has been effectively applied to text generation tasks to enhance the quality of generated responses. For example, the work by An et al. (2022) on contrastive neural text generation demonstrates the benefits of using contrastive objectives to distinguish between similar and dissimilar contexts, thereby improving the coherence and relevance of generated text. Additionally, Liu et al. (2021a) introduced a topic-aware contrastive learning approach for abstractive dialogue summarization. By contrasting different dialogue contexts and responses, their model better captures thematic shifts and important content, leading to more accurate summaries.
3 Methodology
3.1 Task Formalization
Formally, let $D=\{({u_{1}},{s_{1}}),...,({u_{n}},{s_{m}})\}$ represent a multi-party conversation, which contains $n$ utterances and $m$ speakers, and $u_{i}$ is the utterance made by speaker $s_{j}$. Our objective is to develop a model $f$ that can generate a response $r$ for a specific speaker $s_{t}$ based on the context of the dialogue $D$. This can be expressed as: $r=f(D,s_{t})$. In the subsequent sections, we will describe our proposed model, CMR (Contrastive Learning-based Multi-Party Response Generation), which leverages contrastive learning in a two-stage training process to address these components effectively.
3.2 Model Overview
(a) The figure displays the training process of the stage one.
(b) The figure displays the training process of the stage two.
Figure 2: The figure displays the training processes of the CMR model. Stage One (a) shows the encoder training with contrastive learning, where the encoder processes the query and distinguishes between positive and negative samples. Stage Two (b) illustrates the joint training with response generation and contrastive learning, involving both the encoder and decoder. The combined loss of cross-entropy and InfoNCE enhances the model’s ability to generate contextually relevant responses and understand the dialogue context.
CMR is designed to effectively handle the complexities of multi-party dialogue response generation. The model employs a two-stage training process. In the first stage, the encoder is trained with contrastive learning to distinguish between the speaking styles and contexts of different participants. In the second stage, the model undergoes joint training with the response generation task and contrastive learning, enhancing its ability to generate contextually relevant and accurate responses.
3.3 Stage I: Enhancing Encoder Discrimination
The primary focus of stage I is to improve the encoder’s capacity to understand and differentiate between diverse dialogue contexts. Specifically, the encoder undergoes training using contrastive learning by contrasting the utterances of different speakers within the same conversation and those of random speakers from the entire dataset. Thus, the encoder can better distinguish between various speaking styles and related facts, enhancing its ability to capture the unique characteristics of each speaker.
3.3.1 Selection of Positive and Negative Samples
First, we randomly select an utterance from the conversation as the query $q$. The positive sample $k^{+}$ for this query is another utterance spoken by the same speaker within the same conversation. The negative samples ${k^{-}}$ consist of two parts: utterances spoken by other speakers within the same conversation, and utterances spoken by other speakers from the entire dataset, i.e., in-batch samples. It is worth noting that when selecting the query and positive and negative samples, we removed meaningless and dull replies, such as "yes yes yes," to ensure the quality and informativeness of the training data. The detailed methodology for data cleaning is provided in Appendix. By effectively distinguishing between these positive and negative samples, the encoder learns to better capture the nuances of multi-party dialogues.
3.3.2 Training Process of Stage I
The training objective for the first stage is to maximize the similarity between the query and its positive sample while minimizing the similarity between the query and its negative samples. The approach can help the model better understand "who says what" and the stylistic differences between different speakers. We achieve this by using the InfoNCE loss Oord et al. (2018) function.
$\mathcal{L}_{\text{stage-1 }}=-log\frac{exp(\frac{h_{q}\cdot h_{k^{+}}}{\tau})}{{\textstyle\sum_{i=1}^{N}}exp((\frac{h_{q}\cdot h_{k_{i}}}{\tau})+{\textstyle\sum_{j=1}^{T}}exp(\frac{h_{q}\cdot h_{k^{-}_{j}}}{\tau})}$ (1)
$h_{q}=Pooling(Encoder(q))$ (2)
$h_{k^{+}}=Pooling(Encoder(k^{+}))$ (3)
$h_{k^{-}}=Pooling(Encoder(k^{-}))$ (4)
where $h_{q}$ is the representation of the query, $h_{k^{+}}$ is the representation of the positive sample, $h_{k^{-}_{j}}$ represents the negative samples from the same conversation, $h_{k_{i}}$ represents the samples from-batch, and $\tau$ represents temperature. Pooling refers to utilizing the hidden state of the $<$/s$>$ token to represent the encoding of the entire sequence. Because our early experiments revealed that during generation, the decoder’s cross-attention in the T5 model focuses predominantly on the $<$/s$>$ token. Consequently, the $<$/s$>$ token effectively acts as an aggregator of context information, making it a suitable representation of the entire sequence. The parameters $T$ and $N$ denote the number of negative samples from the same conversation and the batch size, respectively.
3.4 Stage II: Joint Training with Response Generation and Contrastive Learning
After the encoder has been trained to distinguish between different speakers and their styles in Stage I, the next step involves joint training with response generation and contrastive learning, which aims to further refine the model’s ability to generate contextually relevant and coherent responses.
3.4.1 Selection of Positive and Negative Samples
As with traditional contrastive learning in generation tasks, we use the dialogue context as the query and the gold response as the positive sample. However, our approach introduces an innovative method for selecting negative samples, which consists of four parts:
  • Previous Utterance by the Target Speaker: The first part of the negative samples is the utterance immediately preceding the gold response spoken by the target speaker. In general, the utterance is highly related to the gold response but contextually different enough due to the progression of the dialogue. We believe the utterance serving as a negative sample can help the model understand the dialogue context by distinguishing between closely related but semantically distinct utterances.
  • Utterances by Other Speakers in the Current Conversation: The second part includes utterances from other speakers within the same conversation with query. These samples help the model grasp the context by understanding topic shifts and interactions between different participants.
  • From Batch Samples: The third part consists of negative samples drawn from the entire dataset (batch samples). These provide a broader context and ensure that the model generalizes well across different dialogues.
  • Model-Generated Beam Search Samples: The fourth part includes the model’s own beam search samples. Following An et al. (2022), these samples, generated by the model itself, are similar in structure to the gold response but differ in content. Including these helps the model refine its understanding and improve the quality of generated responses.
3.4.2 Training Process of Stage II
In Stage Two, the model is trained to generate responses while simultaneously leveraging contrastive learning to enhance the quality of these responses. For the response generation task, we use maximum likelihood estimation with cross-entropy loss. For the contrastive learning part, we use the InfoNCE loss, similar to Stage One. Thus, the training objective in Stage Two combines the response generation loss and the contrastive learning loss. The combined loss function can be expressed as:
$\mathcal{L}_{\text{stage-2 }}=\mathcal{L}_{\text{gen }}+\lambda\mathcal{L}_{\text{cl}}$ (5)
where $\mathcal{L}_{\text{gen }}$ is the response generation loss, calculated using cross-entropy loss between the generated response and the gold response. $\mathcal{L}_{cl}$ is the contrastive learning loss, and $\lambda$ is a hyperparameter that balances the two loss components.
The cross-entropy loss for response generation is defined as:
$\mathcal{L}_{\text{gen }}=-\sum_{t=1}^{T}\log P\left(r^{t}\mid r^{<t},D,s_{t}\right)$ (6)
Here, $r^{t}$ represents the token at position $t$ in the gold response, $r^{<t}$ represents the sequence of tokens in the gold response before position $t$. $D$ is the dialogue context, and $s_{t}$ is the target speaker. The cross-entropy loss measures the negative log-likelihood of the gold response given the dialogue context and the target speaker.
The contrastive learning loss $\mathcal{L}_{cl}$ is defined as:
$\mathcal{L}_{cl}=-log\frac{exp(\frac{h_{q}\cdot h_{k^{+}}}{\tau})}{exp(\frac{h_{q}\cdot h_{k^{+}}}{\tau})+{\textstyle\sum_{j=1}^{P+T+N+B}}exp(\frac{h_{q}\cdot h_{k^{-}_{j}}}{\tau})}$ (7)
where $h_{q}$ is the representation of the query, $h_{k^{+}}$ is the representation of the positive sample, and $h_{k^{-}_{j}}$ represents the negative samples. The representations $h_{q}$ and $h_{k}$ for the query and the samples are obtained in a manner similar to Stage One. The parameters $P,T,N,B$ denote the number of negative samples from four different sources.
The entire training process can be referenced in Appendix 1.
4 Experiments
4.1 Dataset
For evaluating the performance of our proposed CMR model, we considered two widely-used benchmarks in multi-party dialogue research: the Friends datasetShmueli and Ku (2019) and the Ubuntu IRC benchmarkLowe et al. (2015). The Friends dataset, which contains transcripts from the TV show "Friends," features interactions between six main characters with distinct and consistent personalities, providing a rich context for modeling speaking styles. These well-defined personas allow the model to generate dialogue that aligns with each character’s unique traits. In contrast, the Ubuntu IRC benchmark includes conversations from a large number of speakers, many of whom only contribute a single utterance. This makes it challenging to model individual speaking styles. Additionally, the Ubuntu dataset includes addressee tags and other labels that are not typically available in natural multi-party dialogues and differ from the objectives of our model. Given these factors, we determined that the Friends dataset is more suitable for evaluating our model’s ability. After data pre-processing, there are 5000 conversations for training and 653 conversations for testing.
4.2 Comparison Models
Models Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
BartLewis et al. (2020) 4.80 4.78 1.64 0.78 4.46
T5Raffel et al. (2020) 4.92 5.19 1.65 0.67 4.65
HeterMPCGu et al. (2022a) 5.13 5.30 1.41 0.40 4.74
MADNetGu et al. (2023) 6.17 6.00 1.54 0.63 5.95
CONTAn et al. (2022) 6.46 6.19 2.32 1.06 6.18
LLama3-8b-prompt 4.03 4.38 1.07 0.38 3.75
LLama3-8b-ICT 4.91 4.92 1.57 0.79 4.64
LLama3-70b-prompt 6.21 5.94 1.72 0.75 5.58
LLama3-70b-ICT 6.33 6.01 1.70 0.91 5.72
CMR 8.14 8.63 2.95 1.39 7.45
Table 1: The table shows the performance of different models on various automatic evaluation metrics. Our CMR model shows significant improvements over the baseline models. For the specific settings of the contrastive model, please refer to the appendix.
To evaluate the performance of our proposed CMR model, we compared it against several state-of-the-art models for dialogue response generation, including both general-purpose and multi-party dialogue models. Specifically, we considered BART Lewis et al. (2020) and T5 Raffel et al. (2020), both powerful transformer-based models known for their strong performance across NLP tasks, including dialogue generation. We also evaluated HeterMPC Gu et al. (2022a), which uses a heterogeneous graph neural network to capture complex interactions in multi-party dialogues, and MADNet Gu et al. (2023), which applies fully connected graph neural networks to address scenarios with missing addressee labels. Additionally, we included CONT An et al. (2022), which leverages contrastive learning to improve text generation by constructing contrastive examples from the model’s own predictions. Lastly, we used LLama3, a popular large language model with enhanced architecture and training techniques, as a baseline, testing various sizes and employing both prompt-based and in-context learning to assess its performance on our dataset.
4.3 Implementation Details
Models Metrics Rel. Coh. Eng. SA.
Bart 1.42 3.01 1.23 1.25
T5 1.84 2.67 1.66 1.23
HeterMPC 1.46 2.67 1.56 1.52
MADNet 2.12 2.56 2.13 1.87
CONT 2.61 3.21 3.04 2.27
CMR 3.06 3.40 3.09 2.88
Table 2: The table shows the performance of different models on various human evaluation metrics. Our CMR model shows significant improvements over the baseline models.
Our CMR model is implemented using PyTorch, with T5-base serving as the backbone. The learning rate is set to $4e^{-5}$ and we use the AdamW optimizer. The training procedure is divided into two stages. In Stage One, the encoder is trained with contrastive learning for 10 epochs with a batch size of 4. For each batch, 7 negative samples are used, with 4 samples coming from the same conversation and 3 from other samples within the same batch. After completing Stage One, Stage Two involves joint training with response generation and contrastive learning for 2 epochs, also with a batch size of 4. In this stage, 8 negative samples are used, distributed as follows: 1 sample from the previous utterance by the target speaker, 2 samples from other speakers within the same conversation, 3 samples from the same batch, and 2 model-generated beam search samples. The temperature $\tau$ is setting to 0.1. The model training was conducted on an RTX Titan GPU.
4.4 Automatic Evaluation Results
As shown in Table 1, to evaluate the performance of our proposed CMR model, we conducted experiments comparing it against several state-of-the-art models: BART, T5, CONT, HeterMPC, and MADNet. The evaluation metrics used include F1 Score, BLEU, and ROUGE-L, which provide a comprehensive assessment of the quality, relevance, and fluency of the generated responses. what is more, the results are averaged over five runs.
Compared to its backbone model, T5-base, CMR shows substantial improvements across all metrics, with some metrics nearly doubling. In particular, compared to CONT, which is also a contrastive learning-based response generation model, our model also shows significant improvements. Specifically, the F1 Score of CMR is 26.02% higher than that of CONT. For BLEU-1, BLEU-2, BLEU-3 and ROUGE-L, CMR achieves relative improvements of 39.26%, 27.16%, 31.13% and 20.58%, respectively. These results suggest that our approach to integrating contrastive learning with multi-party dialogue response generation is highly effective.
4.5 Human Evaluation Results
In addition to automatic evaluation, we conducted human evaluations (Table 2) to assess the quality of responses generated by our CMR model compared to baseline models. We used four metrics for evaluation: Relevance, Coherence, Engagement, and Speaker Appropriateness. Here, Speaker Appropriateness checks if the response matches the target speaker’s style and persona. Annotators rated each response on a scale of 1 to 5, with dialogue contexts specifically selected to require responses from one of the six main characters in "Friends." Annotators, blind to the model that generated each response and were instructed to reviewing training set utterances to familiarize themselves with speaking style of the main characters.
Models Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
CMR 8.14 8.63 2.95 1.39 7.45
-w/o SS1 7.37 7.64 2.55 1.26 6.77
-w/o SE1 7.83 7.99 2.61 1.31 7.02
-w/o Stage I 7.17 7.04 2.57 1.21 6.86
-w/o PU 7.88 7.89 2.74 1.19 7.09
-w/o SS2 6.93 6.96 2.48 1.07 6.41
-w/o SE2 7.05 6.89 2.68 1.34 6.64
-w/o BS 6.80 6.52 2.58 1.31 6.39
-w/o Stage II 6.41 6.32 2.24 1.04 5.91
Table 3: The table shows the performance of the CMR model and its variants with specific components removed.
As shown in Table 2, our model outperformed the baseline models in all metrics. Notably, general pre-trained models like BART and T5 often produced dull responses, with around 65% of their outputs resembling generic phrases like "I am sorry" or "yes yes yes." This resulted in lower scores for relevance, engagement, and speaker appropriateness, despite relatively high coherence scores due to their logical consistency. Meanwhile, graph-based models like HeterMPC and MADNet struggled without additional annotations, with HeterMPC frequently generating meaningless phrases across all samples. Finally, while graph structures helped improve relevance and coherence, particularly for HeterMPC and MADNet, contrastive learning in models like CMR proved more effective in capturing different speaking styles, demonstrating the unique advantages of each approach.
5 Analysis and Discussion
5.1 Ablation Study
To understand the contribution of each component of our CMR model, we conducted a series of ablation experiments. Each ablation removes a specific component or modifies the model’s training process, allowing us to observe the impact on performance. The results of these ablation studies are summarized in Table 3.
5.1.1 Impact of SS
The removal of SS1 (same conversation’s other speakers’ utterances as negative samples in Stage 1) and SS2 (same conversation’s other speakers’ utterances as negative samples in Stage 2) both resulted in performance drops across all metrics. Specifically, the F1 score decreased by 0.77 points when SS1 was removed and by 1.21 points when SS2 was removed. These results indicate the importance of incorporating context-specific negative samples to enhance the model’s ability to distinguish between relevant and irrelevant utterances.
5.1.2 Impact of Stage I
Removing the entire Stage I resulted in a significant performance drop across all metrics. The F1 score decreased by 0.97 points, BLEU-1 by 1.59 points, BLEU-2 by 0.38 points, BLEU-3 by 0.18 points, and ROUGE-L by 0.59 points. This demonstrates that the contrastive learning stage is crucial for improving the model’s understanding of dialogue context and speaker styles.
$N_{1},T_{1},N_{2},T_{2},P,B$ Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
$1,1,1,1,1,1$ 7.84 8.00 2.63 1.40 7.09
$2,2,2,1,1,1$ 7.95 8.47 2.69 1.37 7.20
$3,3,3,1,1,1$ 7.96 8.44 2.76 1.40 7.19
$\mathbf{4,4,4,2,1,2}$ 8.14 8.63 2.95 1.39 7.45
$5,5,5,3,1,3$ 8.17 8.64 2.87 1.41 7.51
$6,6,6,3,2,3$ 8.20 8.76 2.89 1.37 7.54
$8,8,8,4,3,4$ 8.02 8.58 2.77 1.34 7.20
Table 4: Performance of the CMR model with different numbers of negative samples ($N_{1}$, $T_{1}$, $N_{2}$, $T_{2}$, $P$, $B$). Increasing the number of negative samples initially improves performance, but beyond a certain point, the results start to fluctuate, indicating that excessively high numbers of negative samples can introduce irrelevant or low-quality negatives, which can degrade the model’s effectiveness.
5.1.3 Impact of PU and BS
In Stage 2, removing PU (previous utterance by the target speaker as a negative sample) and BS (beam search generated utterances as negative samples) led to noticeable declines in performance. These results suggest that PU helps the model achieve a fine-grained understanding of the dialogue context, while BS significantly enhances the overall quality of the generated responses.
5.1.4 Impact of Stage II
Removing the entire Stage II resulted in the most significant performance drop across all metrics. The F1 score decreased by 1.73 points, BLEU-1 by 2.31 points, BLEU-2 by 0.71 points, BLEU-3 by 0.35 points, and ROUGE-L by 1.54 points. This underscores the critical role of the joint training phase in fine-tuning the model’s response generation capabilities. Without Stage II, the model struggles to maintain the coherence and relevance of responses, demonstrating the necessity of this stage for achieving high-quality dialogue generation.
5.2 Impact of the Number of Negative Samples
The number of negative samples is crucial for contrastive learning. Therefore, we explored the impact of varying the number of negative samples on the performance of CMR. As shown in Table 4, increasing the number of negative samples initially improves the experimental results. However, beyond a certain point, the performance begins to fluctuate and does not continue to improve steadily. This fluctuation may be due to the inclusion of more irrelevant or less similar negative samples as the number of negatives increases. For instance, as the value of k increases in beam search-generated top-k negative samples, more less optimal negative samples are included, which can negatively impact the effectiveness of contrastive learning.
5.3 Visualization of the Impact of Speaker aware in Stage I
Figure 3: Attention weights of the $<$/s$>$ token in the encoder for (a) the CMR model with Stage I contrastive learning, and (b) the CMR model without Stage I contrastive learning. The attention weights are more focused on contextually relevant tokens such as names and pronouns in (a), indicating improved context understanding after Stage I.
To further understand the impact of Speaker aware in Stage I on the performance of the CMR model, we visualized the representations learned by the model with and without Stage I, as shown in Figure 3. Specifically, we visualized the attention weights of the $<$/s$>$ token to other tokens in the encoder. The reason for choosing the $<$/s$>$ token, as mentioned earlier, is based on our previous experiments. We found that during generation, the cross-attention weights of the decoder in the T5 model are primarily concentrated on the $<$/s$>$ token in the encoder. This indicates that the $<$/s$>$ token acts as an information aggregator in encoder, and its attention weights to other context tokens can effectively reflect the model’s primary focus. As shown in Figure 2, the attention weights of the $<$/s$>$ token in the CMR model with Stage I show a more concentrated pattern. The attention is focused on specific tokens that are contextually relevant, such as names (“Mona", “Ross") and pronouns (“I," “you"), indicating a better understanding of the dialogue context.
6 Conclusion
In this paper, we proposed CMR, a novel contrastive learning-based multi-party dialogue response generation model. Our approach leverages a two-stage training process to enhance the model’s understanding of dialogue context and speaker styles. In Stage One, we employ contrastive learning to differentiate between relevant and irrelevant utterances, significantly improving the model’s ability to comprehend complex multi-party dialogues. Stage Two combines response generation with contrastive learning, further refining the model’s capability to produce contextually relevant and coherent responses. In conclusion, our CMR model represents a significant advancement in multi-party dialogue response generation, offering a robust framework that effectively integrates contrastive learning to tackle the intricate dynamics of multi-party conversations. We believe that our findings will inspire further research and development in this challenging area, ultimately leading to more advanced and capable dialogue systems.
7 Limitations
While the CMR model demonstrates strong performance in multi-party dialogue response generation, several limitations must be acknowledged.
  • The use of contrastive learning, particularly in a multi-stage training process, results in significantly increased training times compared to simpler models. Constructing negative samples and distributed training further adds to the computational cost.
  • Although the model performs well on the Friends dataset, which has clear and well-defined speaker identities, its ability to generalize to other domains or dialogues with less structured or noisier data has not been thoroughly explored.
  • While the model effectively captures speaker-specific characteristics in datasets with distinct characters, such as Friends, it may not perform as well in scenarios where speakers frequently change or where speaker-specific features are less defined. This limits its effectiveness in more spontaneous, open-domain dialogues.
Appendix A The algorithm of entire training process
See Algorithm 1.
Algorithm 1 Training CMR Model
1:Stage One: Encoder Training
2:for epoch in num_epochs_stage_1 do
3:   for batch in dataloader do
4:      Initialize $loss_{stage\_1}\leftarrow 0$
5:      for dialogue, speaker in batch do
6:         Select query $q$ from dialogue
7:         Select positive sample from same speaker
8:         Select negative samples from other speakers and batch
9:         Compute InfoNCE loss
10:         $loss_{stage\_1}\leftarrow loss_{stage\_1}+loss_{InfoNCE}$
11:      end for
12:      Backpropagation and optimization
13:   end for
14:end for
15:Stage Two: Joint Training
16:for epoch in num_epochs_stage_2 do
17:   for batch in dataloader do
18:      Initialize $loss_{gen}\leftarrow 0$, $loss_{cl}\leftarrow 0$
19:      for dialogue D, target speaker $s_{t}$ in batch do
20:         Generate response $\hat{r}$ and negative samples from beam search
21:         Get gold response $r$
22:         Compute generation loss: $loss_{gen}\leftarrow loss_{gen}+\mathcal{L}_{gen}$
23:         Select query $q\leftarrow D$
24:         Select positive sample $r$
25:         Select negative samples from previous utterance, other speakers, batch, beam search
26:         Compute InfoNCE loss: $loss_{cl}\leftarrow loss_{cl}+\mathcal{L}_{cl}$
27:      end for
28:      Combine losses: $loss_{stage\_2}\leftarrow loss_{gen}+\lambda\cdot loss_{cl}$
29:      Backpropagation and optimization
30:   end for
31:end for
Appendix B Clean Friends Dataset
The Friends dataset we obtained included many low-quality samples. Some samples contained numerous meaningless and repetitive phrases, while others were too short and involved only two speakers. Additionally, some samples were fragmented and lacked context, making them unsuitable for training a robust dialogue model.
Bad Words
yes okay hey oh
no hello yeah well
aha ok i am
fine too bye is
good hm alright really
you are ohh uh
huh ha what go
hi thanks right many
and ohhh could can
that
Table 5: List of Bad Words
To address these issues, we first set a list of “bad words," as shown in Table 5. If a sample’s response contained these bad words in more than 30% of its content, the sample was removed. Additionally, we removed samples where the dialogue involved only two speakers. Finally, we manually removed samples that lacked sufficient context.
Appendix C Configuration of Comparison Models
The original HeterMPC and MADNet models use BART-base as their backbone model. We first tested their performance with this configuration. Then, we replaced their backbone model with BART-large. During this modification, we adjusted the “num_gat_heads" parameter in their configuration files to 8, as the original setting of 6 was not divisible by the 1024-dimensional hidden state of BART-large. The results reported in the paper use BART-large as the backbone because both models achieved better performance with this configuration.
For the configurations of LLaMA-3 using Prompt and in-context learning, refer to Table 6. In our experiments, the prompt shown in the table yielded the best results and is used in the reported experiments. The in-context learning configuration involved testing with 1 to 6 samples, and we found that using 3 samples provided the best performance.
Prompt $X_{t}$. You participated in
this conversation as $S_{t}$.
Please provide your
response.
[…]
Table 6: Configurations for LLaMA-3 using Prompt and In-context Learning. $t$ represents the target sample, and $i$ represents a random sample from the training set.

### v2

@@
-Advancing Multi-Party Dialogue Systems with Speaker-ware Contrastive Learning
+Advancing Multi-Party Dialogue Framework with Speaker-ware Contrastive Learning
 Abstract
-Dialogue response generation has made significant progress, but most research has focused on dyadic dialogue.
-In contrast, multi-party dialogues involve more participants, each potentially discussing different topics, making the task more complex.
-Current methods often rely on graph neural networks to model dialogue context, which helps capture the structural dynamics of multi-party conversations.
-However, these methods are heavily dependent on intricate graph structures and dataset annotations, and they often overlook the distinct speaking styles of participants.
-To address these challenges, we propose CMR, a Contrastive learning-based Multi-party dialogue Response generation model.
-CMR uses self-supervised contrastive learning to better distinguish "who says what." Additionally, by comparing speakers within the same conversation, the model captures differences in speaking styles and thematic transitions.
-To the best of our knowledge, this is the first approach to apply contrastive learning in multi-party dialogue generation.
-Experimental results show that CMR significantly outperforms state-of-the-art models in multi-party dialogue response tasks.
-Advancing Multi-Party Dialogue Systems with Speaker-ware Contrastive Learning
+Multi-party dialogues, common in collaborative scenarios like brainstorming sessions and negotiations, pose significant challenges due to their complexity and diverse speaker roles.
+Current methods often use graph neural networks to model dialogue context, capturing structural dynamics but heavily relying on annotated graph structures and overlooking individual speaking styles.
+To address these challenges, we propose CMR, a Contrastive learning-based Multi-party dialogue Response generation framework.
+CMR employs a two-stage self-supervised contrastive learning framework.
+First, it captures global differences in speaking styles across individuals.
+Then, it focuses on intra-conversation comparisons to identify thematic transitions and contextually relevant facts.
+To the best of our knowledge, this is the first approach that applies contrastive learning in multi-party dialogue generation.
+Experimental results demonstrate that CMR not only significantly outperforms state-of-the-art models, but also generalizes well to large pre-trained language models, effectively enhancing their capability in handling multi-party conversations.
+Advancing Multi-Party Dialogue Framework with Speaker-ware Contrastive Learning
 Zhongtian Hu1, Yiwen Cui1, Ronghan Li2, Meng Zhao3, Lifang Wang1, 1School of Computer Science and Engineering, Northwestern Polytechnical University, 2School of Computer Science and Technology, Xidian University, 3School of Artificial Intelligence and Big Data, Henan University of Technology ahxchzt@mail.nwpu.edu.cn, wanglf@nwpu.edu.cn
 1 Introduction
-In recent years, dialogue response generation has made notable progress Ide and Kawahara (2021); Liu et al.
-(2021c), largely due to advancements in language models.
-These improvements have enhanced the generation of coherent and contextually relevant responses in applications such as customer service, virtual assistants, and conversational agents Hu et al.
+Dialogue response generation has made notable progress Ide and Kawahara (2021); Liu et al.
+(2021c) in applications Hu et al.
 (2024); Wu et al.
 (2022); Komeili et al.
-(2022).
+(2022), largely due to advancements in language models.
 However, most research has focused on dyadic dialogues Cai and Cai (2022); Sun et al.
@@
 (2022), involving only two participants.
-In contrast, multi-party dialogues, common in real-world scenarios like group discussions, meetings, and social media, present unique structural challenges.
-As shown in Figure 1, while dyadic dialogues follow natural speaker rotations, multi-party dialogues require careful management of information flow between multiple participants Gu et al.
+In contrast, multi-party dialogues, common in real-world scenarios such as group discussions, meetings, and role-play in game, present unique structural challenges.
+As shown in Figure 1, multi-party dialogues require careful management of information flow between multiple participants Gu et al.
 (2022b); Ganesh et al.
-(2023), posing significant challenges for response generation models.
+(2023).
 Figure 1: This figure illustrates a multi-party dialogue involving four participants: John, Mike, Sarah, and Emma.
@@
 For example, Gu et al.
-(2022a) proposed the HeterMPC model, which uses utterance and interlocutor nodes with six meta relations to construct a heterogeneous graph that captures dialogue interactions.
+(2022a) propose the HeterMPC model, which uses utterance and interlocutor nodes with six meta relations to construct a heterogeneous graph that captures dialogue interactions.
 Similarly, Gu et al.
-(2023) introduced a fully connected graph to model utterance-interlocutor relationships, incorporating latent addressee inference.
-While GNNs effectively model the dependencies in multi-party dialogues, they have drawbacks.
-The complexity of graph structures and the need for extensive annotations, such as addressee labels Li and Zhao (2023), make these methods computationally expensive and challenging to scale in real-world applications.
-Given these challenges, we propose CMR (Contrastive Learning-based Multi-Party Response Generation).
-Our approach uses contrastive learning to understand the structure and dynamics of multi-party dialogues without relying on complex graph structures or extensive annotations.
+(2023) introduce a fully connected graph to model utterance-interlocutor relationships, incorporating latent addressee inference.
+While GNNs effectively model the dependencies in multi-party dialogues, they suffer from four major limitations: (1) they require additional annotations such as speaker interaction labels; (2) they model each dialogue in isolation, failing to capture consistent speaker styles across conversations; (3) they are insensitive to topic transitions, limiting their ability to track discourse dynamics.
+(4) their graph-based architectures are inherently difficult to scale or integrate into large pre-trained language models.
+We propose CMR (Contrastive learning-based Multi-party Response Generation), a novel framework that leverages self-supervised contrastive learning to model the structure and dynamics of multi-party dialogues without relying on complex graph structures or extensive annotations.
 The model follows a two-stage training process.
-In the first stage, only the encoder is trained.
-Through contrastive learning, the model contrasts utterances from different speakers within the same conversation and those from random speakers across the dataset.
-This helps the model distinguish between different facts and speaking styles, enhancing speaker awareness and leading to more relevant responses.
-In the second stage, the model is jointly trained on a response generation task with contrastive learning.
+Stage I focuses on learning to differentiate speakers based on their utterances.
+By contrasting utterances from different speakers within the same conversation and those from random speakers across the dataset against the target speaker’s utterances, the model learns to distinguish speaking styles among different participants.
+Stage II then optimizes the model for response generation with an integrated contrastive objective.
 Here, the model contrasts the dialogue context with generated responses (via beam search), the gold response, the target speaker’s previous utterance, and other speakers’ responses within the same conversation.
-This stage improves response quality by enhancing the model’s understanding of conversation themes.
-We validated our approach using the Friends dataset. Both automatic evaluation results and human evaluation results demonstrate that our model significantly outperforms the baseline models.
+This stage improves response quality by enhancing the model’s understanding of conversation themes and speaker-specific facts.
+We validate our approach on both the Friends dataset and the Ubuntu Dialogue Corpus.
+Specifically, we evaluate CMR on two different backbone models: a traditional encoder-decoder model (T5) and a decoder-only large language model (LLaMA 3.1).
+Evaluation results demonstrate that our model significantly outperforms the baseline models.
 Our primary contributions are as follows:
-  • We propose CMR, a novel multi-party dialogue response generation model that leverages contrastive learning to better understand the structure and dynamics of multi-party dialogues.
-To the best of our knowledge, this is the first work on multi-party dialogue response generation based on contrastive learning.
-  • We introduce a two-stage training process: the first stage focuses on distinguishing the speaking styles and facts among different participants through contrastive learning, and the second stage jointly trains the model with the response generation task and contrastive learning to improve the quality of responses and the model’s understanding of dialogue themes.
-  • Both automatic evaluation results and human evaluation results on the Friends dataset demonstrate that our model significantly outperforms baseline models and achieves SOTA.
+  • We propose CMR, a novel multi-party dialogue response generation framework that leverages contrastive learning to better understand the structure and dynamics of multi-party dialogues.
+  • We introduce a two-stage training process: the first stage focuses on distinguishing the speaking styles, and the second stage improves the quality of responses and the model’s understanding of dialogue themes and speaker-specific facts.
+  • Evaluation results demonstrate that our model significantly outperforms baseline models and achieves SOTA.
 2 Related Work
@@
 2.1 Graph Neural Networks for Multi-Party Dialogue
-Several studies have explored the use of GNNs to capture the intricate relationships in multi-party dialogues.
-For instance, GSN (Graph-Structured Network) introduced by Hu et al.
-(2019).
-models multi-party dialogues as graphs rather than sequences.
-GSN employs an utterance-level graph-structured encoder to model information flow along the graph and incorporates speaker information to improve performance Similarly, Gu et al.
-(2022a) proposed the HeterMPC model, which employs heterogeneous graph networks to model interactions between interlocutors and utterances using multiple types of nodes and edges.
-Furthermore, another approach by Gu et al.
-(2023) constructed fully connected graphs to model relationships between all utterances and interlocutors in a conversation.
-While addressee labels are not entirely unnecessary, their model is designed to infer latent addressees through a graph prediction model integrated into a Variational Inference framework, making it robust even when some addressee information is missing.
+Several studies model multi-party dialogues using graph neural networks For instance, GSN (Graph-Structured Network) Hu et al.
+(2019) uses utterance-level graphs to capture dialogue flow and speaker dynamics.
+Similarly, HeterMPC Gu et al.
+(2022a), which employs heterogeneous graph networks to model interactions between interlocutors and utterances using multiple types of nodes and edges.
+Furthermore, MADNet Gu et al.
+(2023) utilizes fully connected graphs integrated with latent addressee inference via variational methods.
 2.2 Enhanced Speaker-Aware Approaches
 Ma et al.
-(2023) proposed an Enhanced Speaker-Aware Multi-Party Multi-Turn Dialogue Comprehension model that integrates speaker-related features at both the token and discourse levels.
-By applying a speaker mask for attention and using heterogeneous graph networks for comprehensive speaker-aware discourse clues, this model achieves state-of-the-art performance on datasets such as MolweniLi et al.
-(2020) and FriendsQAYang and Choi (2019).
+(2023) introduce a model that applies a speaker mask attention for comprehensive speaker-aware discourse clues.
 Liu et al.
-(2021b) proposed an Utterance-Aware and Speaker-Aware Multi-turn Dialogue Representation model (MDFN) that integrates utterance-level and speaker-level features using a mask attention mechanism.
-The MDFN achieve amazing performance on several Datasets for response selection task.
-These methods highlights the importance of incorporating speaker-aware signals to improve dialogue comprehension and response generation.
+(2021b) propose an Utterance-Aware and Speaker-Aware Multi-turn Dialogue Representation model (MDFN) that integrates utterance-level and speaker-level features using a mask attention mechanism.
+These methods highlight the importance of incorporating speaker-aware signals.
 2.3 Contrastive Learning for Text Generation
@@
 (2021a) introduced a topic-aware contrastive learning approach for abstractive dialogue summarization.
-By contrasting different dialogue contexts and responses, their model better captures thematic shifts and important content, leading to more accurate summaries.
 3 Methodology
@@
 This can be expressed as: $r=f(D,s_{t})$.
-In the subsequent sections, we will describe our proposed model, CMR (Contrastive Learning-based Multi-Party Response Generation), which leverages contrastive learning in a two-stage training process to address these components effectively.
 3.2 Model Overview
-(a) The figure displays the training process of the stage one.
-(b) The figure displays the training process of the stage two.
-Figure 2: The figure displays the training processes of the CMR model.
-Stage One (a) shows the encoder training with contrastive learning, where the encoder processes the query and distinguishes between positive and negative samples.
-Stage Two (b) illustrates the joint training with response generation and contrastive learning, involving both the encoder and decoder.
-The combined loss of cross-entropy and InfoNCE enhances the model’s ability to generate contextually relevant responses and understand the dialogue context.
-CMR is designed to effectively handle the complexities of multi-party dialogue response generation.
-The model employs a two-stage training process.
-In the first stage, the encoder is trained with contrastive learning to distinguish between the speaking styles and contexts of different participants.
-In the second stage, the model undergoes joint training with the response generation task and contrastive learning, enhancing its ability to generate contextually relevant and accurate responses.
-3.3 Stage I: Enhancing Encoder Discrimination
-The primary focus of stage I is to improve the encoder’s capacity to understand and differentiate between diverse dialogue contexts.
-Specifically, the encoder undergoes training using contrastive learning by contrasting the utterances of different speakers within the same conversation and those of random speakers from the entire dataset.
-Thus, the encoder can better distinguish between various speaking styles and related facts, enhancing its ability to capture the unique characteristics of each speaker.
-3.3.1 Selection of Positive and Negative Samples
-First, we randomly select an utterance from the conversation as the query $q$.
-The positive sample $k^{+}$ for this query is another utterance spoken by the same speaker within the same conversation.
-The negative samples ${k^{-}}$ consist of two parts: utterances spoken by other speakers within the same conversation, and utterances spoken by other speakers from the entire dataset, i.e., in-batch samples.
-It is worth noting that when selecting the query and positive and negative samples, we removed meaningless and dull replies, such as "yes yes yes," to ensure the quality and informativeness of the training data.
-The detailed methodology for data cleaning is provided in Appendix.
-By effectively distinguishing between these positive and negative samples, the encoder learns to better capture the nuances of multi-party dialogues.
-3.3.2 Training Process of Stage I
-The training objective for the first stage is to maximize the similarity between the query and its positive sample while minimizing the similarity between the query and its negative samples.
-The approach can help the model better understand "who says what" and the stylistic differences between different speakers.
-We achieve this by using the InfoNCE loss Oord et al.
-(2018) function.
+(a) The figure displays the training process of the Stage I.
+(b) The figure displays the training process of the Stage II.
+Figure 2: The figure displays the training processes of the CMR framework.
+Here, an encoder-decoder architecture is presented as an example.
+Stage I (a) shows the encoder training with contrastive learning.
+Stage II (b) illustrates the joint training with response generation and contrastive learning.
+The framework employs a two-stage training process.
+In the first stage, the model learns utterance-level representations using contrastive learning, enabling it to distinguish between the speaking styles and contexts of different participants.
+In the second stage, the model undergoes joint training with the response generation task and contrastive learning, enhancing its understanding of dialogue themes and speaker-specific facts to generate contextually relevant and accurate responses.
+To provide a detailed understanding of the model’s mechanics, we divide the discussion into two sections: Stage I (Sec 3.3) and Stage II(Sec 3.4).
+Each section covers the selection of positive and negative samples, as well as the training process specific to that stage.
+3.3 Stage I: Speaker Discrimination Learning
+We use contrastive learning to differentiate speaker styles, contrasting utterances within and across conversation.
+3.3.1 Stage I: Selection of Samples
+We select samples for contrastive learning as follows:
+  1. 1.
+    Query ($q$): An utterance randomly selected from a conversation.
+  2. 2.
+    Positive Sample ($k^{+}$): Another utterance spoken by the same speaker within the same conversation.
+  3. 3.
+    Negative Samples (${k^{-}}$):
+    • Utterances spoken by other speakers within the same conversation.
+    • Utterances from random speakers across the dataset (in-batch samples).
+It is worth noting that when selecting the query and positive and negative samples, we remove meaningless and dull replies, such as "yes yes yes," to ensure the quality and informativeness of the training data.
+The detailed methodology for data cleaning is provided in Appendix B.
+3.3.2 Stage I: Training Objective and Process
+We optimize the InfoNCE loss Oord et al.
+(2018) to bring utterances from the same speaker closer while pushing apart those from different speakers, encouraging the model to better understand "who says what" and stylistic differences between different speakers:
 $\mathcal{L}_{\text{stage-1 }}=-log\frac{exp(\frac{h_{q}\cdot h_{k^{+}}}{\tau})}{{\textstyle\sum_{i=1}^{N}}exp((\frac{h_{q}\cdot h_{k_{i}}}{\tau})+{\textstyle\sum_{j=1}^{T}}exp(\frac{h_{q}\cdot h_{k^{-}_{j}}}{\tau})}$ (1)
-$h_{q}=Pooling(Encoder(q))$ (2)
-$h_{k^{+}}=Pooling(Encoder(k^{+}))$ (3)
-$h_{k^{-}}=Pooling(Encoder(k^{-}))$ (4)
-where $h_{q}$ is the representation of the query, $h_{k^{+}}$ is the representation of the positive sample, $h_{k^{-}_{j}}$ represents the negative samples from the same conversation, $h_{k_{i}}$ represents the samples from-batch, and $\tau$ represents temperature.
-Pooling refers to utilizing the hidden state of the $<$/s$>$ token to represent the encoding of the entire sequence.
-Because our early experiments revealed that during generation, the decoder’s cross-attention in the T5 model focuses predominantly on the $<$/s$>$ token.
-Consequently, the $<$/s$>$ token effectively acts as an aggregator of context information, making it a suitable representation of the entire sequence.
-The parameters $T$ and $N$ denote the number of negative samples from the same conversation and the batch size, respectively.
-3.4 Stage II: Joint Training with Response Generation and Contrastive Learning
-After the encoder has been trained to distinguish between different speakers and their styles in Stage I, the next step involves joint training with response generation and contrastive learning, which aims to further refine the model’s ability to generate contextually relevant and coherent responses.
-3.4.1 Selection of Positive and Negative Samples
-As with traditional contrastive learning in generation tasks, we use the dialogue context as the query and the gold response as the positive sample.
-However, our approach introduces an innovative method for selecting negative samples, which consists of four parts:
-  • Previous Utterance by the Target Speaker: The first part of the negative samples is the utterance immediately preceding the gold response spoken by the target speaker.
-In general, the utterance is highly related to the gold response but contextually different enough due to the progression of the dialogue.
-We believe the utterance serving as a negative sample can help the model understand the dialogue context by distinguishing between closely related but semantically distinct utterances.
-  • Utterances by Other Speakers in the Current Conversation: The second part includes utterances from other speakers within the same conversation with query.
-These samples help the model grasp the context by understanding topic shifts and interactions between different participants.
-  • From Batch Samples: The third part consists of negative samples drawn from the entire dataset (batch samples).
-These provide a broader context and ensure that the model generalizes well across different dialogues.
-  • Model-Generated Beam Search Samples: The fourth part includes the model’s own beam search samples.
-Following An et al.
-(2022), these samples, generated by the model itself, are similar in structure to the gold response but differ in content.
-Including these helps the model refine its understanding and improve the quality of generated responses.
-3.4.2 Training Process of Stage II
-In Stage Two, the model is trained to generate responses while simultaneously leveraging contrastive learning to enhance the quality of these responses.
+where the parameters $T$ and $N$ denote the number of negative samples from the same conversation and the batch size, respectively.
+$h_{q}$ represents the embedding of the query utterance, $h_{k^{+}}$ is the embedding of the positive sample, and $h_{k^{-}_{j}}$ are embeddings of negative samples drawn from the same conversation, $h_{k_{i}}$ represents the in-batch samples, and $\tau$ represents the temperature.
+Specifically, for encoder-decoder architectures, these embeddings are obtained by encoding each utterance separately through the encoder and applying a pooling operation over the encoder’s token-level hidden states.
+Thus, only encoder are trained for encoder-decoder model in Stage I.
+For decoder-only architectures, the embeddings are obtained by passing each utterance through the transformer-decoder individually and applying mean-pooling across the token-level hidden states produced by the decoder.
+3.4 Stage II: Enhancing Contextual Response
+This stage involves joint training with response generation and contrastive learning, which aims to further refine the model’s ability to generate contextually relevant and coherent responses.
+3.4.1 Stage II: Selection of Samples
+For a specific conversation, we select samples of the following types:
+  1. 1.
+    Query ($q$): Dialogue context.
+  2. 2.
+    Positive Sample ($k^{+}$): The gold response for the target speaker.
+  3. 3.
+    Negative Samples ($k^{-}$): The negative samples combine those from Stage I with additional types:
+    • The target speaker’s previous utterance.
+    • Model-generated beam search samples.
+3.4.2 Stage II: Training Objective and Process
 For the response generation task, we use maximum likelihood estimation with cross-entropy loss.
-For the contrastive learning part, we use the InfoNCE loss, similar to Stage One.
-Thus, the training objective in Stage Two combines the response generation loss and the contrastive learning loss.
+For the contrastive learning part, we use the InfoNCE loss, similar to Stage I.
+Thus, the training objective in Stage II combines the response generation loss and the contrastive learning loss.
 The combined loss function can be expressed as:
@@
 $\mathcal{L}_{cl}=-log\frac{exp(\frac{h_{q}\cdot h_{k^{+}}}{\tau})}{exp(\frac{h_{q}\cdot h_{k^{+}}}{\tau})+{\textstyle\sum_{j=1}^{P+T+N+B}}exp(\frac{h_{q}\cdot h_{k^{-}_{j}}}{\tau})}$ (4)
-where $h_{q}$ is the representation of the query, $h_{k^{+}}$ is the representation of the positive sample, and $h_{k^{-}_{j}}$ represents the negative samples.
-The representations $h_{q}$ and $h_{k}$ for the query and the samples are obtained in a manner similar to Stage One.
+The representations $h_{q}$ and $h_{k}$ for the query and samples are obtained in a manner similar to Stage I.
 The parameters $P,T,N,B$ denote the number of negative samples from four different sources.
-The entire training process can be referenced in Appendix 1.
+The entire training process can be referenced in Appendix A.
 4 Experiments
 4.1 Dataset
-For evaluating the performance of our proposed CMR model, we considered two widely-used benchmarks in multi-party dialogue research: the Friends datasetShmueli and Ku (2019) and the Ubuntu IRC benchmarkLowe et al.
+To evaluate the performance of CMR, we consider two widely-used benchmarks in multi-party dialogue research: the Friends datasetShmueli and Ku (2019) and the Ubuntu IRC benchmarkLowe et al.
 (2015).
 The Friends dataset, which contains transcripts from the TV show "Friends," features interactions between six main characters with distinct and consistent personalities, providing a rich context for modeling speaking styles.
-These well-defined personas allow the model to generate dialogue that aligns with each character’s unique traits.
-In contrast, the Ubuntu IRC benchmark includes conversations from a large number of speakers, many of whom only contribute a single utterance.
-This makes it challenging to model individual speaking styles.
-Additionally, the Ubuntu dataset includes addressee tags and other labels that are not typically available in natural multi-party dialogues and differ from the objectives of our model.
-Given these factors, we determined that the Friends dataset is more suitable for evaluating our model’s ability.
-After data pre-processing, there are 5000 conversations for training and 653 conversations for testing.
+In contrast, the Ubuntu dataset includes many speakers with sparse utterances, posing challenges for style modeling.
+Given these factors, we determine that the Friends dataset is more suitable for evaluating our model’s ability.
+However, to assess the generalization capability of our model, we apply the second-stage trained model to the Ubuntu dataset.
 4.2 Comparison Models
 Models Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
+Friends
 BartLewis et al. (2020) 4.80 4.78 1.64 0.78 4.46
@@
 LLaMA3-8b-ICT 4.91 4.92 1.57 0.79 4.64
+LLaMA3-8b-lora 5.37 5.26 1.87 0.80 4.91
 LLaMA3-70b-prompt 6.21 5.94 1.72 0.75 5.58
 LLaMA3-70b-ICT 6.33 6.01 1.70 0.91 5.72
-CMR 8.14 8.63 2.95 1.39 7.45
+LLaMA3-70b-lora 5.33 5.12 2.11 0.98 4.90
+CMR-T5 8.14 8.63 2.95 1.39 7.45
+CMR-LLaMA3 8b 8.71 9.13 3.16 1.44 7.82
+Ubuntu
+GSN*Hu et al. (2019) - 9.02 2.78 - -
+ASRG*Song et al. (2022) - 9.41 2.89 - -
+ChatMDG*Li et al. (2024) - 9.72 3.04 - -
+BartLewis et al. (2020) 11.71 11.07 3.75 1.62 10.35
+T5Raffel et al. (2020) 12.04 11.91 3.65 1.74 10.69
+HeterMPCGu et al. (2022a) 11.95 11.58 3.87 1.71 10.46
+MADNetGu et al. (2023) 12.13 11.64 3.81 1.78 11.13
+LLaMA3-8b-lora 12.75 12.39 3.98 1.80 12.07
+LLaMA3-70b-lora 12.61 12.43 4.03 1.66 12.00
+CMR-T5(only Stage II) 12.41 12.37 3.99 1.80 11.89
+CMR-LLaMA3 8b (only Stage II) 13.43 13.01 4.28 1.86 12.89
 Table 1: The table shows the performance of different models on various automatic evaluation metrics.
-Our CMR model shows significant improvements over the baseline models.
-For the specific settings of the contrastive model, please refer to the appendix.
-To evaluate the performance of our proposed CMR model, we compared it against several state-of-the-art models for dialogue response generation, including both general-purpose and multi-party dialogue models.
-Specifically, we considered BART Lewis et al.
-(2020) and T5 Raffel et al.
-(2020), both powerful transformer-based models known for their strong performance across NLP tasks, including dialogue generation.
-We also evaluated HeterMPC Gu et al.
-(2022a), which uses a heterogeneous graph neural network to capture complex interactions in multi-party dialogues, and MADNet Gu et al.
-(2023), which applies fully connected graph neural networks to address scenarios with missing addressee labels.
-Additionally, we included CONT An et al.
-(2022), which leverages contrastive learning to improve text generation by constructing contrastive examples from the model’s own predictions.
-Lastly, we used LLama3, a popular large language model with enhanced architecture and training techniques, as a baseline, testing various sizes and employing both prompt-based and in-context learning to assess its performance on our dataset.
+For the specific settings of the LLaMA3 model, please refer to the appendix C.
+* indicates that we directly report the results obtained from training on the original UBUNTU dataset as reported in the original paper.
+To evaluate the performance of CMR, we compare it against several state-of-the-art models for dialogue response generation, including both general-purpose and multi-party dialogue models.
+Evaluated models include graph-based methods (GSNHu et al.
+(2019), HeterMPC Gu et al.
+(2022a), MADNet Gu et al.
+(2023), ChatMDGLi et al.
+(2024)), contrastive learning methods (CONT An et al.
+(2022)), and transformer-based models (BART Lewis et al.
+(2020), T5 Raffel et al.
+(2020)).
+Notably, BART and T5 also serve as the backbone for all the aforementioned architectures.
+Lastly, we evaluate LLaMA3.1, a popular large language model.
 4.3 Implementation Details
+Our CMR framework is combined with different backbone models, T5 and LLaMA 3.1.
+For LLaMA-based models, we apply LoRA during training to enable parameter-efficient fine-tuning.
+The learning rate is set to $4e^{-5}$ and we use the AdamW optimizer.
+In Stage I, the model is trained with contrastive learning for 10 epochs with a batch size of 4.
+For each batch, 7 negative samples are used, with 4 samples coming from the same conversation and 3 from other samples within the same batch.
+Stage II involves joint training with response generation and contrastive learning for 2 epochs, also with a batch size of 4 and $\lambda$ is setting to 2.
+In this stage, 8 negative samples are used, distributed as follows: 1 sample from the previous utterance by the target speaker, 2 samples from other speakers within the same conversation, 3 samples from the same batch, and 2 model-generated beam search samples.
+The temperature $\tau$ is set to 0.1.
+4.4 Automatic Evaluation Results
+As shown in Table 1, to evaluate the performance of our proposed CMR framework, we conduct experiments comparing it against several state-of-the-art models. The results are averaged over five runs.
+On the Friends dataset, CMR demonstrates significant improvements across all evaluation metrics when combined with different backbone models, with several metrics nearly doubling.
+Specifically, compared to CONT, another contrastive learning-based dialogue response generation model, CMR-T5 substantially increases the F1 score by 26.02%, and achieves notable gains on BLEU and ROUGE scores.
+Additionally, compared to graph-based models such as HeterMPC and MADNet, CMR-T5 also exhibits substantial performance improvements.
+Meanwhile, the decoder-only LLM variant, CMR-LLaMA, significantly outperforms the vanilla LLaMA model, further highlighting the effectiveness and generalizability of our approach.
+For the Ubuntu dataset, due to its data format being incompatible with the full CMR model, we apply only CMR (Stage II) for evaluation.
+Despite this limitation, CMR still outperform all other models of the same scale.
+4.5 LLM-Based Pairwise Preference Evaluation
+We observe that for such complex multi-party dialogue scenarios, conventional automatic evaluation metrics may be insufficient to fully capture the quality differences between model responses.
+To address this, we employ the GPT-4o model as an LLM-based judge, instructing it to carefully consider aspects such as speaker style consistency, personalization, contextual coherence, and response informativeness.
+Specifically, we randomly sample 500 dialogue instances from the Friends dataset and conduct pairwise preference comparisons among responses generated by CMR-T5, CMR-LLaMA, LLaMA 3.1 8B, and LLaMA 3.1 70B.
+As shown in Figure 3, although CMR-LLaMA achieves only slightly better scores than CMR-T5 on automatic metrics, the LLM-based preference evaluation shows that CMR-LLaMA is preferred in 66% of the comparisons, indicating its superior ability to generate high-quality responses in multi-party dialogue settings.
+4.6 Human Evaluation Results
+In addition to automatic evaluation, we conduct human evaluations (Table 2) to assess the quality of responses generated by our CMR-T5 model compared to baseline models.
+We use four metrics for evaluation: Relevance, Coherence, Engagement, and Speaker Appropriateness.
+Here, Speaker Appropriateness checks if the response matches the target speaker’s style and persona.
+Annotators rate each response on a scale of 1 to 5, with dialogue contexts specifically selected to require responses from one of the six main characters in "Friends." Annotators, who are blind to the model that generated each response and were instructed to reviewing training set utterances to familiarize themselves with speaking style of the main characters.
+Figure 3: LLM-based pairwise preference comparison on the Friends dataset.
 Models Metrics Rel. Coh. Eng. SA.
@@
 MADNet 2.12 2.56 2.13 1.87
-CONT 2.61 3.21 3.04 2.27
+CONT-T5 2.61 3.21 3.04 2.27
 CMR 3.06 3.40 3.09 2.88
 Table 2: The table shows the performance of different models on various human evaluation metrics. Our CMR model shows significant improvements over the baseline models.
-Our CMR model is implemented using PyTorch, with T5-base serving as the backbone.
-The learning rate is set to $4e^{-5}$ and we use the AdamW optimizer.
-The training procedure is divided into two stages.
-In Stage One, the encoder is trained with contrastive learning for 10 epochs with a batch size of 4.
-For each batch, 7 negative samples are used, with 4 samples coming from the same conversation and 3 from other samples within the same batch.
-After completing Stage One, Stage Two involves joint training with response generation and contrastive learning for 2 epochs, also with a batch size of 4.
-In this stage, 8 negative samples are used, distributed as follows: 1 sample from the previous utterance by the target speaker, 2 samples from other speakers within the same conversation, 3 samples from the same batch, and 2 model-generated beam search samples.
-The temperature $\tau$ is setting to 0.1.
-The model training was conducted on an RTX Titan GPU.
-4.4 Automatic Evaluation Results
-As shown in Table 1, to evaluate the performance of our proposed CMR model, we conducted experiments comparing it against several state-of-the-art models: BART, T5, CONT, HeterMPC, and MADNet.
-The evaluation metrics used include F1 Score, BLEU, and ROUGE-L, which provide a comprehensive assessment of the quality, relevance, and fluency of the generated responses.
-what is more, the results are averaged over five runs.
-Compared to its backbone model, T5-base, CMR shows substantial improvements across all metrics, with some metrics nearly doubling.
-In particular, compared to CONT, which is also a contrastive learning-based response generation model, our model also shows significant improvements.
-Specifically, the F1 Score of CMR is 26.02% higher than that of CONT.
-For BLEU-1, BLEU-2, BLEU-3 and ROUGE-L, CMR achieves relative improvements of 39.26%, 27.16%, 31.13% and 20.58%, respectively.
-These results suggest that our approach to integrating contrastive learning with multi-party dialogue response generation is highly effective.
-4.5 Human Evaluation Results
-In addition to automatic evaluation, we conducted human evaluations (Table 2) to assess the quality of responses generated by our CMR model compared to baseline models.
-We used four metrics for evaluation: Relevance, Coherence, Engagement, and Speaker Appropriateness.
-Here, Speaker Appropriateness checks if the response matches the target speaker’s style and persona.
-Annotators rated each response on a scale of 1 to 5, with dialogue contexts specifically selected to require responses from one of the six main characters in "Friends." Annotators, blind to the model that generated each response and were instructed to reviewing training set utterances to familiarize themselves with speaking style of the main characters.
 Models Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
@@
 Table 3: The table shows the performance of the CMR model and its variants with specific components removed.
-As shown in Table 2, our model outperformed the baseline models in all metrics.
-Notably, general pre-trained models like BART and T5 often produced dull responses, with around 65% of their outputs resembling generic phrases like "I am sorry" or "yes yes yes." This resulted in lower scores for relevance, engagement, and speaker appropriateness, despite relatively high coherence scores due to their logical consistency.
-Meanwhile, graph-based models like HeterMPC and MADNet struggled without additional annotations, with HeterMPC frequently generating meaningless phrases across all samples.
-Finally, while graph structures helped improve relevance and coherence, particularly for HeterMPC and MADNet, contrastive learning in models like CMR proved more effective in capturing different speaking styles, demonstrating the unique advantages of each approach.
+As shown in Table 2, our model outperforms the baseline models in all metrics.
+Notably, general pre-trained models like BART and T5 often produce dull responses, with around 65% of their outputs resembling generic phrases like "I am sorry" or "yes yes yes." This results in lower scores for relevance, engagement, and speaker appropriateness, despite relatively high coherence scores due to their logical consistency.
+Meanwhile, graph-based models like HeterMPC and MADNet struggle without additional annotations, with HeterMPC frequently generating meaningless phrases across all samples.
+Finally, while graph structures help improve relevance and coherence, particularly for HeterMPC and MADNet, contrastive learning in models like CMR proves more effective in capturing different speaking styles, demonstrating the unique advantages of each approach.
 5 Analysis and Discussion
+All analyses in the following sections are conducted based on the CMR-T5 model.
 5.1 Ablation Study
-To understand the contribution of each component of our CMR model, we conducted a series of ablation experiments.
+To understand the contribution of each component of our CMR model, we conduct a series of ablation experiments.
 Each ablation removes a specific component or modifies the model’s training process, allowing us to observe the impact on performance.
+As mentioned earlier, we conducted ablation experiments on Friends to isolate component impacts, as Ubuntu’s format limits full CMR application.
 The results of these ablation studies are summarized in Table 3.
@@
 5.1.2 Impact of Stage I
-Removing the entire Stage I resulted in a significant performance drop across all metrics.
+Removing the entire Stage I results in a significant performance drop across all metrics.
 The F1 score decreased by 0.97 points, BLEU-1 by 1.59 points, BLEU-2 by 0.38 points, BLEU-3 by 0.18 points, and ROUGE-L by 0.59 points.
 This demonstrates that the contrastive learning stage is crucial for improving the model’s understanding of dialogue context and speaker styles.
-$N_{1},T_{1},N_{2},T_{2},P,B$ Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
-$1,1,1,1,1,1$ 7.84 8.00 2.63 1.40 7.09
-$2,2,2,1,1,1$ 7.95 8.47 2.69 1.37 7.20
-$3,3,3,1,1,1$ 7.96 8.44 2.76 1.40 7.19
-$\mathbf{4,4,4,2,1,2}$ 8.14 8.63 2.95 1.39 7.45
-$5,5,5,3,1,3$ 8.17 8.64 2.87 1.41 7.51
-$6,6,6,3,2,3$ 8.20 8.76 2.89 1.37 7.54
-$8,8,8,4,3,4$ 8.02 8.58 2.77 1.34 7.20
-Table 4: Performance of the CMR model with different numbers of negative samples ($N_{1}$, $T_{1}$, $N_{2}$, $T_{2}$, $P$, $B$).
-Increasing the number of negative samples initially improves performance, but beyond a certain point, the results start to fluctuate, indicating that excessively high numbers of negative samples can introduce irrelevant or low-quality negatives, which can degrade the model’s effectiveness.
 5.1.3 Impact of PU and BS
@@
 5.2 Impact of the Number of Negative Samples
-The number of negative samples is crucial for contrastive learning.
-Therefore, we explored the impact of varying the number of negative samples on the performance of CMR.
-As shown in Table 4, increasing the number of negative samples initially improves the experimental results.
-However, beyond a certain point, the performance begins to fluctuate and does not continue to improve steadily.
-This fluctuation may be due to the inclusion of more irrelevant or less similar negative samples as the number of negatives increases.
-For instance, as the value of k increases in beam search-generated top-k negative samples, more less optimal negative samples are included, which can negatively impact the effectiveness of contrastive learning.
+The number of negative samples is crucial for contrastive learning. Therefore, we explored the impact of varying the number of negative samples on the performance of CMR. As shown in Appendix D.
 5.3 Visualization of the Impact of Speaker aware in Stage I
-Figure 3: Attention weights of the $<$/s$>$ token in the encoder for (a) the CMR model with Stage I contrastive learning, and (b) the CMR model without Stage I contrastive learning.
-The attention weights are more focused on contextually relevant tokens such as names and pronouns in (a), indicating improved context understanding after Stage I.
-To further understand the impact of Speaker aware in Stage I on the performance of the CMR model, we visualized the representations learned by the model with and without Stage I, as shown in Figure 3.
-Specifically, we visualized the attention weights of the $<$/s$>$ token to other tokens in the encoder.
-The reason for choosing the $<$/s$>$ token, as mentioned earlier, is based on our previous experiments.
-We found that during generation, the cross-attention weights of the decoder in the T5 model are primarily concentrated on the $<$/s$>$ token in the encoder.
-This indicates that the $<$/s$>$ token acts as an information aggregator in encoder, and its attention weights to other context tokens can effectively reflect the model’s primary focus.
-As shown in Figure 2, the attention weights of the $<$/s$>$ token in the CMR model with Stage I show a more concentrated pattern.
-The attention is focused on specific tokens that are contextually relevant, such as names (“Mona", “Ross") and pronouns (“I," “you"), indicating a better understanding of the dialogue context.
+To further understand the impact of Speaker aware in Stage I on the performance of the CMR-T5 model, we visualize the representations learned by the model with and without Stage I, as shown in Appendix E.
+5.4 Generalization and Application Scenarios.
+CMR is particularly well-suited for scenarios where a fixed set of speakers appears across dialogues, such as NPC conversations in games or episodic scripted content.
+In these cases, modeling consistent speaker styles across samples can bring notable benefits.
+Nonetheless, CMR is not restricted to such settings.
+On the Ubuntu dataset where speakers rarely repeat, CMR still outperforms all baseline models of the same scale, despite being slightly behind LLM.
+We can say that the second stage of our framework alone is effective in modeling multi-party dialogues, even in the absence of speaker-level recurrence.
+Moreover, unlike prior methods, CMR requires no additional annotations, further demonstrating its adaptability to diverse dialogue scenarios.
+To further evaluate the generalization ability of CMR, we conduct additional experiments in settings where the model encounters speakers with limited presence in the training data.
+As detailed in the appendix F, CMR still achieves strong performance under this condition.
+Intuitively, the phenomenon may be attributed to the model’s ability to generalize from frequently seen speakers.
+While the model may be unable to fully capture the speaking style of rarely seen speakers, it can effectively leverage its understanding of other speakers to distinguish between interlocutors.
+Thus, CMR remains capable of identifying speaker-specific factual content and maintaining consistency in speaker-related information.
+5.5 Isolating Speaker Style: Comparison with Per-Speaker Models
+One of the core motivations of CMR is to model individualized speaker style within multi-party dialogues.
+A natural question arises: would training a separate model for each speaker yield better style modeling, as each model can specialize in the language patterns of a single character?
+To address this question, we conduct an experiment on the Friends dataset by training six separate models, one for each of the six main characters.
+During inference, we invoke the model corresponding to the ground-truth speaker.
+Please note that although the Friends dataset includes additional speakers beyond the six main ones, this experiment is restricted to training and testing only on the six main characters.
+As such, the results of CMR in this setting are slightly different from those reported earlier.
+The details are shown in Appendix G, the experimental results clearly demonstrate that training separate models for each speakers significantly poorer performance compared to CMR.
+5.6 Robustness to Varying Speaker Counts
+To evaluate the robustness of our model in handling dialogues with different levels of speaker complexity, we conducted additional experiments on subsets of dialogues containing 3, 4, 5, and 6 speakers.
+Using the Entire Dataset.
+We first tested our model, trained on the entire Friends dataset, on conversations with varying speaker counts.
+We observe a gradual performance decrease as the number of speakers increases.
+However, the decline may be attributed not only to the inherent complexity of multi-party dialogue but also to the training data distribution, which predominantly features conversations with 3 or 4 speakers.
+Using Balanced Subsets.
+To isolate the effect of speaker count, we created balanced subsets of 400 samples for each speaker setting.
+We retrained the model on these balanced subsets.
+Although the overall scores are lower compared to the model trained on the full dataset, the performance decline across speaker counts is notably smaller, suggesting that our model retains a reasonable degree of robustness when faced with increasing dialogue complexity.
+All details and results are shown in Appendix H.
+5.7 Training Cost and Deployment
+Due to CMR’s two-stage training strategy and the need to generate responses for contrastive learning during the second stage, CMR requires more training time than its backbone model.
+However, once training is complete, CMR demonstrates a distinct advantage over graph-based models during deployment.
+The contrastive learning component of CMR is entirely absent during inference, meaning that its resource requirements and inference speed are identical to those of a standard T5 model.
+In contrast, graph-based models require the construction of graphs during both training and inference, which incurs additional computational overhead.
 6 Conclusion
-In this paper, we proposed CMR, a novel contrastive learning-based multi-party dialogue response generation model.
+We propose CMR, a novel contrastive learning-based multi-party dialogue response generation framework.
 Our approach leverages a two-stage training process to enhance the model’s understanding of dialogue context and speaker styles.
-In Stage One, we employ contrastive learning to differentiate between relevant and irrelevant utterances, significantly improving the model’s ability to comprehend complex multi-party dialogues.
-Stage Two combines response generation with contrastive learning, further refining the model’s capability to produce contextually relevant and coherent responses.
-In conclusion, our CMR model represents a significant advancement in multi-party dialogue response generation, offering a robust framework that effectively integrates contrastive learning to tackle the intricate dynamics of multi-party conversations.
-We believe that our findings will inspire further research and development in this challenging area, ultimately leading to more advanced and capable dialogue systems.
+In Stage I, we employ contrastive learning to differentiate between relevant and irrelevant utterances, significantly improving the model’s ability to comprehend complex multi-party dialogues.
+Stage II combines response generation with contrastive learning, further refining the model’s capability to produce contextually relevant and coherent responses.
+In conclusion.
+Our CMR framework advances multi-party dialogue response generation by integrating contrastive learning to effectively handle complex dialogue dynamics and speaker variability, providing a robust foundation for future research in this domain.
 7 Limitations
 While the CMR model demonstrates strong performance in multi-party dialogue response generation, several limitations must be acknowledged.
-  • The use of contrastive learning, particularly in a multi-stage training process, results in significantly increased training times compared to simpler models.
-Constructing negative samples and distributed training further adds to the computational cost.
-  • Although the model performs well on the Friends dataset, which has clear and well-defined speaker identities, its ability to generalize to other domains or dialogues with less structured or noisier data has not been thoroughly explored.
-  • While the model effectively captures speaker-specific characteristics in datasets with distinct characters, such as Friends, it may not perform as well in scenarios where speakers frequently change or where speaker-specific features are less defined.
-This limits its effectiveness in more spontaneous, open-domain dialogues.
+  • While the model does not require additional resources compared to the backbone model during deployment, its training process is time-consuming due to the multi-stage contrastive learning framework.
+  • While the model effectively captures speaker-specific characteristics in datasets with distinct characters, such as Friends, and also demonstrates promising results on Ubuntu by leveraging only Stage II training, it may still be inherently better suited for scenarios where the set of speakers is relatively fixed and speaker-specific features are explicitly available.
 Appendix A The algorithm of entire training process
@@
 Algorithm 1 Training CMR Model
-1:Stage One: Encoder Training
+1:Stage I: Encoder Training
 2:for epoch in num_epochs_stage_1 do
@@
 14:end for
-15:Stage Two: Joint Training
+15:Stage II: Joint Training
 16:for epoch in num_epochs_stage_2 do
@@
 31:end for
-Appendix B Clean Friends Dataset
-The Friends dataset we obtained included many low-quality samples.
+Appendix B Clean Dataset
+The Friends and Ubuntu dataset we obtained included many low-quality samples.
 Some samples contained numerous meaningless and repetitive phrases, while others were too short and involved only two speakers.
@@
 Table 4: List of Bad Words
-To address these issues, we first set a list of “bad words," as shown in Table 5.
-If a sample’s response contained these bad words in more than 30% of its content, the sample was removed.
-Additionally, we removed samples where the dialogue involved only two speakers.
-Finally, we manually removed samples that lacked sufficient context.
-Appendix C Configuration of Comparison Models
-The original HeterMPC and MADNet models use BART-base as their backbone model.
-We first tested their performance with this configuration.
-Then, we replaced their backbone model with BART-large.
-During this modification, we adjusted the “num_gat_heads" parameter in their configuration files to 8, as the original setting of 6 was not divisible by the 1024-dimensional hidden state of BART-large.
-The results reported in the paper use BART-large as the backbone because both models achieved better performance with this configuration.
-For the configurations of LLaMA-3 using Prompt and in-context learning, refer to Table 6.
-In our experiments, the prompt shown in the table yielded the best results and is used in the reported experiments.
-The in-context learning configuration involved testing with 1 to 6 samples, and we found that using 3 samples provided the best performance.
+$N_{1},T_{1},N_{2},T_{2},P,B$ Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
+$1,1,1,1,1,1$ 7.84 8.00 2.63 1.40 7.09
+$2,2,2,1,1,1$ 7.95 8.47 2.69 1.37 7.20
+$3,3,3,1,1,1$ 7.96 8.44 2.76 1.40 7.19
+$\mathbf{4,4,4,2,1,2}$ 8.14 8.63 2.95 1.39 7.45
+$5,5,5,3,1,3$ 8.17 8.64 2.87 1.41 7.51
+$6,6,6,3,2,3$ 8.20 8.76 2.89 1.37 7.54
+$8,8,8,4,3,4$ 8.02 8.58 2.77 1.34 7.20
+Table 5: Performance of the CMR model with different numbers of negative samples ($N_{1}$, $T_{1}$, $N_{2}$, $T_{2}$, $P$, $B$).
+Increasing the number of negative samples initially improves performance, but beyond a certain point, the results start to fluctuate, indicating that excessively high numbers of negative samples can introduce irrelevant or low-quality negatives, which can degrade the model’s effectiveness.
 Prompt $X_{t}$. You participated in
@@
 Table 6: Configurations for LLaMA-3 using Prompt and In-context Learning. $t$ represents the target sample, and $i$ represents a random sample from the training set.
+To address these issues, we first set a list of “bad words," as shown in Table 4.
+If a sample’s response contained these bad words in more than 30% of its content, the sample was removed.
+Additionally, we removed samples where the dialogue involved only two speakers.
+Finally, we manually removed samples that lacked sufficient context.
+After data pre-processing, there are 5000 conversations for training and 653 conversations for testing on Friends dataset.
+For the Ubuntu dataset, since each utterance is relatively short, we also require that each dialogue contains at least 6 utterances and the entire conversation needed to exceed 512 tokens.
+There are about 67000 conversations for Ubuntu dataset.
+Appendix C Configuration of LLaMA
+We use LLaMA 3.1 as our comparison model.
+For the configurations of LLaMA-3 using Prompt and in-context learning, refer to Table 6.
+In our experiments, the prompt shown in the table yielded the best results and is used in the reported experiments.
+The in-context learning configuration involved testing with 1 to 6 samples, and we found that using 3 samples provided the best performance.
+Appendix D Impact of the Number of Negative Samples
+As shown in Table 5, increasing the number of negative samples initially improves the experimental results.
+However, beyond a certain point, the performance begins to fluctuate and does not continue to improve steadily.
+This fluctuation may be due to the inclusion of more irrelevant or less similar negative samples as the number of negatives increases.
+For instance, as the value of k increases in beam search-generated top-k negative samples, more less optimal negative samples are included, which can negatively impact the effectiveness of contrastive learning.
+Appendix E Visualization of the Impact of Speaker aware in Stage I
+Figure 4: Attention weights of the $<$/s$>$ token in the encoder for (a) the CMR model with Stage I contrastive learning, and (b) the CMR model without Stage I contrastive learning.
+The attention weights are more focused on contextually relevant tokens such as names and pronouns in (a), indicating improved context understanding after Stage I.
+As shown in Figure 4, we visualize the attention weights of the $<$/s$>$ token to other tokens in the encoder.
+The reason for choosing the $<$/s$>$ token, as mentioned earlier, is based on our previous experiments.
+We found that during generation, the cross-attention weights of the decoder in the T5 model are primarily concentrated on the $<$/s$>$ token in the encoder.
+This indicates that the $<$/s$>$ token acts as an information aggregator in encoder, and its attention weights to other context tokens can effectively reflect the model’s primary focus.
+As shown in Figure 2, the attention weights of the $<$/s$>$ token in the CMR model with Stage I show a more concentrated pattern.
+The attention is focused on specific tokens that are contextually relevant, such as names (“Mona", “Ross") and pronouns (“I," “you"), indicating a better understanding of the dialogue context.
+Appendix F Generalization to Rare and Unseen Speakers
+To investigate the generalization capabilities of CMR with respect to rarely observed or completely unseen speakers during training, we extract dialogues from the Friends dataset involving speakers beyond the six main characters.
+Specifically, we isolate response samples from the test set where the speakers are infrequently encountered or entirely absent in the training phase.
+These samples are evaluated using the fully trained CMR model without any further fine-tuning and the results in Table 7 illustrate the performance of CMR on responses by rare or unseen speakers.
+Evaluation F1 BLEU-1 BLEU-2 BLEU-3 ROUGE-L
+Rare Speakers 8.11 8.65 2.85 1.38 7.29
+Standard 8.14 8.63 2.95 1.39 7.45
+Table 7: CMR Performance on responses from rare and unseen speakers.
+The experimental outcomes demonstrate that the CMR model maintains strong performance in generating responses even for speakers rarely or never encountered during training.
+A plausible explanation for this generalization ability is that although the CMR model may not fully capture the speaking style of rarely seen or unseen speakers due to limited data exposure, it can effectively leverage the comprehensive understanding gained from frequent speakers.
+With the assistance of the first stage, while the model may be unable to fully capture the speaking style of rarely seen speakers, it can effectively leverage its understanding of other speakers to distinguish between interlocutors.
+The second stage, is less affected by unseen speakers because it primarily focuses on dialogue themes, context continuity, and factual consistency.
+Consequently, CMR remains capable of recognizing thematic shifts and accurately maintaining speaker-related factual content, thereby ensuring coherent and contextually relevant responses despite encountering rarely or never-before-seen speakers.
+Appendix G Training separate models for each speakers
+We extract samples for the six main characters from the Friends dataset and train six baseline models separately.
+During the testing phase, a different baseline model will be invoked according to the speaker to evaluate its performance.
+Method F1 Score BLEU-1 BLEU-2 BLEU-3 ROUGE-L
+Separate models 7.33 7.90 2.61 1.25 7.04
+CMR 8.14 8.63 2.95 1.39 7.45
+Table 8: Comparison between separate models and our proposed method.
+The experimental results clearly demonstrate significantly poorer performance compared to CMR. We attribute this outcome to two primary factors:
+  • The approach of training separate models for each character only enables the model to capture superficial personalized information about individual characters.
+More importantly, it fails to help the model understand the nuanced differences between different characters.
+  • The original Friends dataset has a limited number of samples to begin with. After splitting the training process, each model only has access to approximately one-sixth of the original data.
+Appendix H Robustness to Varying Speaker Counts
+In this section, we provide additional details and full results for our robustness analysis on varying speaker counts.
+Experimental Setup.
+We conducted two sets of experiments to assess the model’s performance when handling conversations with different numbers of participants (3, 4, 5, and 6 speakers):
+  • Experiment 1: Using the Entire Dataset.
+We evaluated the model trained on the full dataset without controlling for speaker count distribution.
+This reflects a realistic but potentially biased setting, as most training samples involve only 3 or 4 speakers.
+  • Experiment 2: Using Balanced Subsets.
+To control for data imbalance, we created balanced subsets with 400 samples each for conversations involving 3, 4, 5, and 6 speakers.
+We retrained the model on each subset and evaluated its performance.
+Results on the Entire Dataset.
+# Speakers F1 BLEU-1 BLEU-2 BLEU-3 ROUGE-L
+3 8.50 9.00 3.10 1.50 7.80
+4 8.21 9.13 2.99 1.47 7.49
+5 7.91 8.47 2.88 1.35 7.36
+6 7.65 8.22 2.73 1.30 7.10
+Table 9: Performance on conversations with 3–6 speakers using the entire dataset.
+Results on Balanced Subsets.
+# Speakers F1 BLEU-1 BLEU-2 BLEU-3 ROUGE-L
+3 7.84 8.46 2.93 1.48 7.28
+4 7.69 8.23 2.82 1.41 7.12
+5 7.69 8.13 2.77 1.39 7.09
+6 7.63 8.22 2.73 1.37 7.11
+Table 10: Performance on balanced subsets with 3–6 speakers (400 samples each).
+Analysis.
+The results across both experimental settings reveal a consistent performance drop as the number of speakers increases.
+However, when using balanced subsets, the performance degradation is much less pronounced, suggesting that our model can generalize reasonably well across varying